OpenStudy (anonymous):

all lines whose slopes and x intecept equal Ans. y=y'x-y'^2

OpenStudy (anonymous):

help

OpenStudy (amoodarya):

what you have tried ?

OpenStudy (anonymous):

\[y=mx+b\]

OpenStudy (anonymous):

y=bx+b y'=b

OpenStudy (anonymous):

subst. y=y'x+y' factor out

OpenStudy (anonymous):

y=y'(x+1)

OpenStudy (anonymous):

what is the next step ? @amoodarya

ganeshie8 (ganeshie8):

Looks you're wrongly assuming \(b\) is the x intercept

ganeshie8 (ganeshie8):

For a straight line, \(y = mx + b\), the x intercept can be worked by plugging in \(y=0\) : \[0 = mx+b \implies x = -\frac{b}{m}\]

ganeshie8 (ganeshie8):

therefore \(-\frac{b}{m}\) is the x intercept, not \(b\)

ganeshie8 (ganeshie8):

Since you want that equal to the slope, \(-\frac{b}{m}=m \implies b=-m^2\) and the equation of line becomes : \[y = mx-m^2\] plugin \(m=y'\) and you're done!

OpenStudy (anonymous):

@ganeshie8 thank you , i did it :)

OpenStudy (anonymous):

@ganeshie8 same answer :)